OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-2, -5) and (2, 3)

OpenStudy (anonymous):

@hartnn @aaronq @Nurali @Destinymasha

OpenStudy (anonymous):

Point slope form is given by: \[y - y_1 = m(x-x_1)\]

OpenStudy (nfcfox):

First, fine da slope

OpenStudy (nfcfox):

can u do that m80

OpenStudy (anonymous):

Find the slope first.. :)

OpenStudy (anonymous):

\[Slope = \frac{y_2 -y_2}{x_2-x_1}\]

OpenStudy (nfcfox):

this easy

OpenStudy (anonymous):

(-2)-2 over -(5)-3 ?

OpenStudy (anonymous):

Oh wait, @nfcfox mind what you are saying..

OpenStudy (nfcfox):

now we use da pointslope

OpenStudy (anonymous):

Both of you, use good language..

OpenStudy (anonymous):

True.

OpenStudy (anonymous):

Again find the slope properly.. :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Use that formula, plug in the values and show me, don't worry if you go wrong, I am here to correct you only.. :)

OpenStudy (anonymous):

-5 - 3 OVER 2-(-2)

OpenStudy (anonymous):

See, you are writing -5 first, then in the denominator, take -2 first..

OpenStudy (anonymous):

-5 - 3 pver -2 - 2

OpenStudy (anonymous):